OpenStudy (anonymous):

1. If the height h of a triangle is 3 inches less than the length of the base b, and the area A of the triangle is 19 times the length of the base, find b and h.

OpenStudy (anonymous):

eq1. h=(b-3) eq2. A=19b since A= bh/2. then A=b(b-3)/2 so: b(b-3)/2=19b simplify: b=38+3 b=41 and h=38